OpenStudy (anonymous):

Use the following four coordinates to determine which one has a distance of 10 units from point C (1,2). (3,4) (7,10) (6,7) (8,9)

OpenStudy (anonymous):

@math92130 @Mokeira @HDStorm @laurisve @amriju @juanpabloJR @j.youmans @Dr.Tehuti-mes @Pretty_in_pink @trexnapora someone please help

OpenStudy (sidsiddhartha):

use the distance formula\[d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\]

OpenStudy (amriju):

Please elaborate what x1, x2..etc are for her help @sidsiddhartha

OpenStudy (mokeira):

@sidsiddhartha do you mean this \[10=\sqrt{(1-x)^{2}+(2-y)^{2}}\] square both sides \[100=1+x ^{2}+4+y ^{2}\] \[x ^{2}+y ^{2}=95\]

OpenStudy (sidsiddhartha):

yeah if there are two points in a plane with coordinates (x1,y1) and (x2,y2) then distence between them is \[d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\] now take \[x_1\rightarrow 1 ~~and ~~~y_1\rightarrow 2\] then it will be \[10=\sqrt{(1-x_2)^2+(2-y_2)^2}\] now squaring both sides \[100=(1-x_2)^2+(2-y_2)^2\]

OpenStudy (anonymous):

I'm confused on what to do next

OpenStudy (sidsiddhartha):

now check which pair satisfies this equation

OpenStudy (anonymous):

This one! (7,10)

OpenStudy (sidsiddhartha):

yup correct :)

OpenStudy (anonymous):

Thank you can you help me with another one

OpenStudy (sidsiddhartha):

yeah

OpenStudy (anonymous):

Line AB has an equation of a line y=5x-2 and line DC has and equation of a line y=-5x-2. These two lines are parallel because the slope of the lines are the same perpendicular because the product of the slopes is -1 both representations of the same line neither parallel nor perpendicular

OpenStudy (anonymous):

I know its not the 1st or 2nd

OpenStudy (sidsiddhartha):

can u find the slopes of those lines

OpenStudy (anonymous):

5x and -5x

OpenStudy (anonymous):

they both have the same y intercept but I don't know how that answers the question

OpenStudy (anonymous):

I think its the last one, but I'm not sure

OpenStudy (sidsiddhartha):

sorryy connection got lost :(

OpenStudy (sidsiddhartha):

yeah now try to compare with the standard equation \[y=mx+c\] comparing u'll get slope for the first line is m1=5 and for the second line m2=-5 so product of m1*m2=-25 which is not equal to -1 so they are not perpendicular and also \[m_1 \neq m_2\] they are also not parallel

OpenStudy (anonymous):

exactly but I don't know wether its the last answer or second to last answer

OpenStudy (anonymous):

although i'm pretty sure its the last one